OpenStudy (yamyam70):

I need help with this! or someone check my work! THanks!!! A right circular cylinder is inside a right circular cone of altitude h and radius base x. A. Derive a formula for the radius r of the cylinder in terms of h and x if its lateral area is equal to the lateral area of the small cone which surmounts the cylinder.

OpenStudy (yamyam70):



OpenStudy (yamyam70):

\[A1/A2 = 2 \pi r h / \pi r \sqrt{h^2+ x^2} \]

OpenStudy (perl):

Assumptions: radius of cylinder is r, radius of cone is x, height of cone= height of cylinder lateral area of cylinder = lateral area of cone This gives us 2Pi*r*h = Pi * x sqrt( h^2 + x^2) Solve for r r = Pi * x sqrt( h^2 + x^2) / (2Pi*h) r = x sqrt(h^2 +x^2) / (2h )

OpenStudy (yamyam70):

Thanks a lot , you really helped me :) Sorry didnt give credit on the right time :)

OpenStudy (perl):

no problem :)

OpenStudy (yamyam70):

:)